OpenStudy (anonymous):

For the simple harmonic motion equation 5sin(2pit), what is the period?

OpenStudy (badhi):

$$A\sin(\omega t)$$ is the general expression Period T is $$T=\frac{2\pi}{\omega}$$ find $$\omega$$ and substitute to find the T

OpenStudy (anonymous):

Thank You @BAdhi

OpenStudy (badhi):

you're welcome :)